Find the future value, using the future value formula and a calculator. (Give your answer to the nearest cent.)
$42.61 at 6.5% compounded semiannually for 6 years

Answer by rfer(12657)

FV=42.61(1+0.065/2)^2*6
FV=42.61*(1.0.325)^12
FV=42.61*1.46785
FV=$62.55